DIRECT MEASUREMENTS OF COLLISIONALLY BROADENED (CO_{2}-CO_{2}) S-BRANCH RAMAN CO-HERENCE LIFETIMES OF CO_{2}

\underline{JOSEPH R. GORD}, Department of Chemistry, Purdue University, West Lafayette, IN 47907; SUKESH ROY, PAUL S. HSU, NAIBO JIANG, WARUNA D. KULATILAKA, and HANS U. STAUFFER, Spectral Energies, LLC, 5100 Springfield Street, Suite 301, Dayton, OH 45431; JAMES R. GORD, Air Force Research Laboratory, Propulsion Directorate, Wright-Patterson AFB, OH 45433.

We report direct measurement of S-branch Raman coherence lifetimes of CO_{2} due to CO_{2}-CO_{2} collisions by employing picosecond time-resolved coherent anti-Stokes Raman scattering (CARS) spectroscopy. A custom-built, high-peak-power, nearly transform-limited ps laser system offers an ideal combination of frequency and temporal resolution for such measurements. The rotational S-branch transitions of CO_{2} ground state [0,0^{0},0] with rotational quantum number J= 0 − 52 were simultaneously excited by using a broadband (∼3-nm) laser pulse with a full-width-half-maximum (FWHM) of ∼100 ps. The coherence lifetimes of self-broadened CO_{2} for a pressure range of 0.05-1.5 bar were directly measured by probing the rotational coherence with a nearly transform-limited, 80-ps-long laser pulse. The measured linewidth of J=6 and J=50 transitions are found to be ∼0.106±0.0002 and ∼0.070±0.0002, respectively. As expected, the energy-transfer from high J levels has a significantly longer coherence lifetime because of the inertia associated with higher angular momentum. These measurements are very significant for performing accurate thermometry or CO_{2} concentration measurements in gas-phase reacting flows.